OpenStudy (anonymous):

What is the sum of the coefficients of ( [3x - 3x^2 +1]^744 ) x ( [- 3x + 3x^2 +1]^745 ) ??

OpenStudy (anonymous):

tough one. i'll give it a go.

OpenStudy (anonymous):

\[\left(1+3 x-3 x^2\right)^{1489} \]

OpenStudy (anonymous):

^ nope... the second term is different to the first term, so you don't just add the indices

OpenStudy (anonymous):

=(-(3x^2-3x-1))^744x(3x^2-3x+1)^745 dont forget about (x+y)(x-y)=(x^2-y^2) =(3x^2-3x+1)((3x^2-3x)^2-1)^744

OpenStudy (anonymous):

Let a = 3x - 3x^2 = (a + 1)^744 * (-a + 1)^745 = (1 + a)^744 * (1 - a)^644 * (1 - a) = (1 - a^2)^744 * (1 - a) = (a^2 - 1)^744 * (1 - a) Now let (a^2 - 1)^744 = M: and then i'll spare you the details, but i get the coefficients of M can be described as: sum(0,1488) 1488Ck * (-1)^k * sum(0,744) 3^(1488-2k) * 744Ck * (-1)^k is there a simpler way ? This is so messy.

OpenStudy (anonymous):

^ nope... the second term is different to the first term, so you don't just add the indices

OpenStudy (anonymous):

Yes, an error.

myininaya (myininaya):

yucky problem lol

OpenStudy (anonymous):

The answer is 1.\[\left(1+3 x-3 x^2\right) \left(1-3 x+3 x^2\right)^2= \]\[1-3 x-6 x^2+45 x^3-90 x^4+81 x^5-27 x^6 \]If x is set to 1, the sum of the coefficients is 1.\[\left(1+3 x-3 x^2\right)^{744} \left(1-3 x+3 x^2\right)^{745}\text{/. }x\to 1\text{ }\text{ is}\text{ }1 \]Mathematica Home Edition was used for the calculations.

OpenStudy (anonymous):

i think almost all of the middle terms will actually cancel out...let me double check..

OpenStudy (anonymous):

An expanded Mathematica solution with comments is attached.